OpenStudy (anonymous):

g(x)=(2+x)(7+x)(8+x) find g'(x)

OpenStudy (anonymous):

how would i do a problem like this

OpenStudy (anonymous):

saiberz, u there?

OpenStudy (anonymous):

(f(x)*g(x))' = f'(x)*g(x) + f(x)*g'(x) so you can do it twice: g(x) = (2+x)'*(7+x)(8+x) + (2+x)*[(7+x)(8+x)]' g(x) = (1)(7+x)(8+x) + (2+x) [(7+x)'(8+x)+(7+x)(8+x)'] g(x) = (7+x)(8+x) + (2+x)*[(1)(8+x)+(7+x)(1)] i hope it helps you

OpenStudy (anonymous):

o so i gotta use the product which i havent mastered yet lol

OpenStudy (anonymous):

i meant to say product rule lol

OpenStudy (anonymous):

so what would be the final answer?

OpenStudy (anonymous):

just operate it g(x) = (7+x)(8+x) + (2+x)*[(1)(8+x)+(7+x)(1)] g(x) = x^2+15x+ 56 + 2x^2 + 19x +30 g(x) = 3x^2 + 34x +86 check it i'm not sure

OpenStudy (anonymous):

o

OpenStudy (anonymous):

2x^2 + 19x +30 howd u get that ^^^

OpenStudy (anonymous):

(2+x)*[(1)(8+x)+(7+x)(1)] is not that?

OpenStudy (anonymous):

can u show me the steps please. sorry if im being complicated. i know how to FOIL but i always forget what to do when theres 3 rather 2 sets to FOIL out.

OpenStudy (anonymous):

u left :(

OpenStudy (anonymous):

just multiply : (2+x)*(8+x + 7 + x) (2+x)*(15+2x) 30 + 15x + 4x +2x^2 2x^2 + 19x + 30

OpenStudy (anonymous):

i add? ...o i see

OpenStudy (anonymous):

Yes, it is correct. Thanks for all the help. It is appreciated.

OpenStudy (anonymous):

you are welcome